OpenStudy (anonymous):

what 2 numbers multiply to get 720 but add to get -54 ???

OpenStudy (lgbasallote):

x + y = -54 xy = 720 therefore x = 720/y now sub 720/y + y = -54 720 + y^2 = -54 y y^2 + 54y + 720 = 0 factor

OpenStudy (anonymous):

i have to solve for x: 6x^2-54x+120=0